OpenStudy (anonymous):

What is the a value of the following function? 1/3 1/2 3 2

OpenStudy (anonymous):

I don't understand it.

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

Thanks. How did you get that?

jimthompson5910 (jim_thompson5910):

hint: solve y = ax^2 for 'a' when x = 1, y = 2 (notice how the point (1,2) lies on the parabola)

OpenStudy (anonymous):

Oh ok, how does that mean it is 1/2?

jimthompson5910 (jim_thompson5910):

it's not 1/2

jimthompson5910 (jim_thompson5910):

that person probably just guessed

OpenStudy (anonymous):

Oh thank you. How do I figure out the answer?

jimthompson5910 (jim_thompson5910):

y = ax^2 2 = a*1^2 a = ???

OpenStudy (anonymous):

What is a?

OpenStudy (anonymous):

I don't understand the equation

jimthompson5910 (jim_thompson5910):

y = x^2 is the parent function of any quadratic

jimthompson5910 (jim_thompson5910):

y = ax^2 vertically stretches or compresses the parabola y = x^2

jimthompson5910 (jim_thompson5910):

the 'a' determines how stretched it is

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

would it be 1/3?

jimthompson5910 (jim_thompson5910):

you have to solve the equation 2 = a*1^2 since 1^2 is the same as 1, you solve 2 = a*1

jimthompson5910 (jim_thompson5910):

have a look at this page http://www.purplemath.com/modules/grphquad2.htm

OpenStudy (anonymous):

Ohh 2

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Thank you :)

jimthompson5910 (jim_thompson5910):

yw